In Anasville nominal GDP in 2030 is $30,000 and the price index (base of 100 in 2020) stands at 150. Find real GDP expressed in 2020 dollars.


1
Expert's answer
2021-09-09T11:05:25-0400

Nominal GDP in 2030 is $30,000

 the price index (base of 100 in 2020) stands at 150

"Real \space GDP =\frac{ Nominal \space GDP}{Price\space Index\space of \space current \space year}\times Price\space Index \space of\space base\space year\\Real\space GDP =\frac{ 30,000}{150}\times100\\Real\space GDP = 200\times 100\\Real\space GDP = 20,000"